OpenStudy (anonymous):

Solve and Check your answer... 5/1-x = 6/1+x

OpenStudy (dls):

\[\LARGE \frac{5}{1-x}=\frac{6}{1+x}\]

OpenStudy (dls):

Cross Multiply

OpenStudy (anonymous):

5(1+x) = 6(1-x)

OpenStudy (anonymous):

x = 1/11

OpenStudy (dls):

\[\LARGE \frac{5}{1-\frac{1}{11}}=\frac{6}{1+\frac{1}{11}}\] Check.

OpenStudy (dls):

If LHS=RHS. Then its correct.